# Feedback-Free Resource Scheduling: Towards Flexible Multi-BS Cooperation in FD-RAN

### v1

@@ V-A Problem Transformation
We modify the original decision variable to enable the transformation of the original optimization problem into an ILP problem. Specifically, we define $\mathcal{M}^{q}$ as the collection of cooperative sets, each consisting of $q$ BSs selected from the BS set $\mathcal{M}$. For example, when there are three BSs in the network, i.e., =$\mathcal{M}=\{1,2,3\}$, the set of all combinations of two BSs selected from $\mathcal{M}$ is $\mathcal{M}^{2}=\{\{1,2\},\{1,3\},\{2,3\}\}$. Next, we introduce a binary variable $\delta_{\mathcal{B},n}$ to represent the selection of cooperation set $\mathcal{B}$ for user $n$. When $\delta_{\mathcal{B},n}=1$, it indicates that cooperation set $\mathcal{B}\in\mathcal{M}^{q}$ is selected to serve user $n$, where $q\in\mathcal{M}$ and $|\mathcal{B}|=q$; if $\delta_{\mathcal{B},n}=0$, it means the cooperation set is not chosen. Let $o_{\mathcal{B},n}$ denote the number of subcarriers allocated to user $n$ by cooperation set $\mathcal{B}$. The constraint that the total number of allocated subcarriers does not exceed the available total number of subcarriers can be transformed into
@@
The objective function in (23) is expressed in terms of the transformed integer decision variables, while constraint (26) ensures that user demands are satisfied. The objective function and all constraints are linear, with decision variables restricted to integers, resulting in a standard ILP problem.
@@ V-B Two-Stage Resource Allocation Algorithm
To solve the aforementioned ILP problem, the TSRA algorithm is proposed. First, the feasibility pump algorithm quickly provides a feasible solution for subcarrier allocation. Then, the branch-and-cut method is applied to iteratively refine the solution to achieve optimality.
@@
$d(\bm{\chi},[\bm{\chi}^{\star}])=\sum_{\begin{subarray}{c}j\in\mathcal{J}\\ [\delta_{j}^{\star}]=1\end{subarray}}(1-\delta_{j})+\sum_{\begin{subarray}{c}j\in\mathcal{J}\\ [\delta_{j}^{\star}]=0\end{subarray}}\delta_{j}$ (29)
$+\sum_{\begin{subarray}{c}j\in\mathcal{J}\\ [o_{j}^{\star}]=K\end{subarray}}(K-o_{j})+\sum_{\begin{subarray}{c}j\in\mathcal{J}\\ [o_{j}^{\star}]=0\end{subarray}}o_{j}+\sum_{\begin{subarray}{c}j\in\mathcal{J}\\ 0<[o_{j}^{\star}]<K\end{subarray}}(o_{j}^{+}+o_{j}^{-}),$
where $o_{j}^{+}$ and $o_{j}^{-}$ represent the non-negative differences between $o_{j}$ and $[o_{j}]$, corresponding to cases where $o_{j}>[o_{j}]$ and $o_{j}<[o_{j}]$, respectively. Then, the formulated LP problem in feasibility pump $\mathcal{Z}_{LP}(\bm{\chi},[\bm{\chi}^{\star}])$ can be expressed as follows with additional constraints on $o_{j}$:
@@
If $d(\bm{\chi},[\bm{\chi}^{\star}])=0$, a feasible solution to the original problem is obtained. The feasibility pump algorithm, forming the first stage of the proposed resource allocation method, is detailed in the initial part of Algorithm 1. Note that in Step 1 of Algorithm 1, $N_{flip}$ elements of $\bm{\chi}^{*}$ are replaced with $\bm{\chi}^{\star}$ rounded in the opposite direction to prevent an endless loop.
@@ 11                   $\bm{\chi}^{*}\leftarrow[\bm{\chi}^{\star}]$
14                   Select $N_{flip}$ variables of $\bm{\chi}^{*}$ with the highest $|\chi_{j}^{\star}-\chi_{j}^{*}|$ and replace $\chi_{j}^{*}$ with $\chi_{j}^{\star}$ rounded in the opposite direction
@@
17            Solve $\mathcal{Z}_{LP}(\bm{\chi}^{\star},\bm{\chi}^{*})$ of (30) to obtain $\bm{\chi}^{\star}$
@@ 44                   Add them to the LP relaxation of the problem $\mathcal{Z}_{y}$ and go back to Step 1
The latter part of Algorithm 1, starting from Step 1, outlines the branch-and-cut method. It utilizes a tree structure to represent the problem-solving process, where the root node corresponds to the LP relaxation of the original problem, providing an initial LB of the subcarrier energy cost. Similarly, the relaxed LP problem is solved by the interior-point method. The LB is the smallest energy cost from all explored LP relaxations, while the UB is the best energy cost from feasible integer solutions, initialized using the results of the feasibility pump algorithm. Through branching, it partitions the feasible regions into smaller subspaces, creating subproblems as child nodes. At each node, cutting planes are introduced to tighten the solution space further, eliminating relaxed solutions that violate the integer constraints. Pruning is applied to discard branches where the LB exceeds the incumbent UB, reducing unnecessary searches. Fig. 3 provides a clear depiction of the workflow of the proposed TSRA algorithm, highlighting the processes of branching, cutting, and pruning.
Figure 3: Illustration of TSRA algorithm using an example with two decision variables. The black solid circles represent feasible integer solutions, while the blue region depicts the solution space under the LP relaxation. The yellow circle indicates the feasible integer solution obtained from the feasibility pump algorithm, with its corresponding subcarrier energy cost initialized as the UB. The red solid square marks the incumbent best solution in the LP relaxation. The purple dashed line represents the added cutting plane.
By integrating the branch-and-cut with the feasibility pump, a feasible solution is quickly obtained and iteratively refined to optimality. The complexity of the proposed TSRA algorithm mainly stems from its second stage, which is the focus of the analysis. Essentially, the branch-and-cut method establishes a search tree to explore and obtain the optimal solution. At each node, the LP relaxation of the problem is efficiently solved by the interior-point method. In the worst case, the time complexity is $\mathcal{O}(2^{|\mathcal{J}|})$ in terms of the number of nodes generated by a brute-force search. However, in practice, the number of actually explored nodes is generally much smaller due to effective pruning and cutting strategies. Moreover, the high-quality feasible solution from the first stage provides an initial UB that further facilitates pruning. Thus, the time complexity of the second stage is $\mathcal{O}(N_{en})$ in terms of the number of actually explored nodes $N_{en}$, where $N_{en}\ll 2^{|\mathcal{J}|}$. The space complexity is $\mathcal{O}(N_{an})$, where $N_{an}$ denotes the number of active nodes during the search, including those being processed and queued for exploration. Effective pruning typically ensures that $N_{an}<N_{en}$.
@@ VI-A Simulation Setup
In our simulation, BS is equipped with $N_{tx}=16$ transmit antennas, and the user is equipped with $N_{rx}=4$ receive antennas. Zero forcing (ZF) equalizer is employed on the user side. The center frequency is 3.5 GHz and the subcarrier spacing is 15 kHz, with a transmission time interval (TTI) of 1 ms and the numerology $\mu=0$. In a single TTI or time slot, the resource element (RE) grid for MIMO transmission consists of 2016 REs, obtained from 144 subcarriers $\times$ 14 symbols, i.e., $K=144$. For simplicity, $OH$ is fixed at 0.14. The transmit power per subcarrier $p$ is set to 1 mW. In the greedy algorithm, $\eta_{\min}$ is fixed at 0.1. In TSRA, $N_{flip}$ is chosen to be 15 to avoid an endless loop and the large constant $\lambda$ is set to 1,000.
@@
Meanwhile, simulations are conducted based on the realistic ray-tracing channel dataset: DeepMIMO [37]. Using accurate ray-tracing data obtained from Remcom Wireless InSite, clustered delay line (CDL) channels between BS and user are generated according to standard [38]. DeepMIMO dataset provides candidate BSs and users with various coordinates, with BSs at a height of 6 m and users at 2 m. In this paper, we focus on using BS1,2,3 to transmit data to users in User Grid 1 of DeepMIMO’s O1 scenario, i.e., $M=3$. A total of 225 users are considered, spanning from R515 to R915 in O1. Among them, 10 users are randomly selected in each case to have data rate demands and are served, which means $N=10$. When generating the channel, the maximum number of channel paths is set to 10, though the actual number varies at different locations based on ray-tracing results.
We perform simulations for both consecutive and independent scheduling decisions. In consecutive scheduling, we simulate time-varying channels by generating 100 distinct channel samples over a 100 ms time period. Specifically, we assume that the channel path loss remains constant at a specific location. As the O1 scenario is static, we adapt it by using ray-tracing results in the CDL model to generate channel data with random variations for each TTI. The random seeds are set from 1 to 100 to reflect the small-scale environmental variations, which yields a channel coherence time of 1 ms. Fig. 4 provides detailed user and BS locations. User data rate demands are determined by the data accumulated in the radio link control (RLC) buffer, calculated as the difference between the total arrived and transmitted data. Given the relatively short simulated time period, the data arriving at each user’s buffer in each TTI is assumed to remain relatively stable. Its size follows a Gaussian distribution with a variance of 0.01 times the mean, indicating the stability of the arriving data. The mean size of arrived data for different users is also modeled as a Gaussian distribution, with a mean of $\rho$ and a variance of 0.2 times the mean, reflecting variations in user services. Moreover, considering the short simulation duration, user locations are assumed to remain constant. In independent scheduling, each test case is a single time slot without data accumulation in the buffer. To evaluate the impact of feedback delay, each test case is chosen as the second time slot from the previously generated 100 samples, enabling comparisons of feedback-based approaches under no feedback delay and 1 ms delay, denoted as "without delay" and "with delay", respectively. Further details are discussed below.
Figure 4: Locations of selected users and BSs in DeepMIMO’s O1 scenario.
Efficient running time is crucial for the practical implementation of scheduling algorithms. Considering the effectiveness of NNs and the widespread adoption of RL in addressing resource scheduling problems [27, 39, 40], the proposed algorithms are benchmarked against RL. Additionally, we compare feedback-free resource allocation with its feedback-based counterpart using the same scheduling algorithm. Based on the amount of feedback data, CSI feedback can be categorized into implicit and explicit feedback. The former provides a precoding matrix indicator (PMI) from a predefined codebook, while the latter reports the optimal precoder. Accordingly, these are referred to as "feedback PMI" and "feedback optimal precoder", respectively. Considering the increasing overheads of pilots and CSI feedback and computational complexity of joint transmission parameters for multi-BS cooperation, feedback-based multi-BS JT is restricted to two BSs, referred to as CoMP. In contrast, FD-RAN enables flexible cooperation among all BSs through the established $\mathcal{FF}_{\mathcal{TM}}$. Unless specified, single BS and CoMP in this paper refer to the transmissions from BS1 and BS1,3, respectively.
VI-B Comparison on Independent Scheduling Decisions
We first verify the effectiveness of the proposed resource scheduling algorithm and compare feedback-free resource scheduling with the feedback-based approach on independent time slots. In these cases, the evaluation assumes an ideal setup with no time delay in running the scheduling algorithm and activating the scheduling decisions, i.e., $d_{2}=d_{3}=0$. Then, the user demand satisfaction rate is calculated as the ratio of transmitted data without error in a time slot to the data arriving within the same time slot. To compare the proposed algorithm with RL under heavy-load and light-load conditions, we generate 500 environments per condition, with random initialization of user locations, data arrivals at the RLC buffer, and user priority weights. We randomly select 80% of the samples for training and use the remaining 20% for testing.
Given that the decision variable is discrete, we adopt the dueling double deep Q-network (D3QN) algorithm as the RL approach to solve the problem. Compared with DQN, D3QN enhances training stability, improves learning efficiency, and results in more accurate Q-value predictions [41, 42]. In this paper, fully connected NNs are utilized in D3QN, consisting of two hidden layers with 128 neurons each. The RL agent undergoes training for 10,000 episodes to ensure convergence. To balance exploration and exploitation, we adopt the $\varepsilon$-decay-greedy strategy. Initially, the probability of random action selection $\varepsilon$ is set to 0.99 and decreases by $10^{-4}$ with each learning step until it reaches 0.01. After 4000 exploration steps, the current Q-network is updated every 4 steps, and the target Q-network every 400 steps. The minibatch size is 64, and ADAM [43] optimizer with $10^{-4}$ learning rate is used for updating neurons’ weights. The activation function is rectified linear unit (ReLU). The memory buffer size is set to 1,000,000.
VI-B1 Evaluation of Proposed Greedy Algorithm
We first compare the proposed greedy algorithm with RL in heavy-load network conditions. For RL, it is essential to specify problem-specific state $\mathfrak{s}$, action $\mathfrak{a}$, reward $\mathfrak{r}$, and episode. Thus, state is defined as the concatenation of each user’s state, i.e., $\mathfrak{s}=(\mathfrak{s}_{1},\ldots,\mathfrak{s}_{N})$, where the state of user $n$ is expressed as $\mathfrak{s}_{n}=(\bm{loc}_{n},\vartheta_{n},R_{n}^{d},\hat{R}_{n})$. Each variable is normalized before being fed into the neural network. Action $\mathfrak{a}$ is designed as allocating one subcarrier from all DL-BSs to a selected user. Therefore, the action space is given by $\mathcal{A}=\{1,\ldots,N\}$, where $\mathfrak{a}\in\mathcal{A}$. Action mask is applied to filter available actions where the user’s demand is unmet, ensuring the generation of meaningful memories for RL to learn effectively [39]. The reward is the increase in the chosen user’s weighted demand satisfaction rate, i.e., $\mathfrak{r}=\vartheta_{\mathfrak{a}}$. At each step of an episode, an action is chosen, transitioning the network to the next state and generating a reward. The episode terminates when all $K_{\text{remaining}}$ subcarriers are allocated. Through interactions with the environment, the RL agent is expected to learn a policy that maximizes the weighted sum of user demand satisfaction rates $V$ based on the received rewards. To demonstrate the superiority of the proposed greedy algorithm over RL, we present the running time and cumulative density function (CDF) of $V$ across 100 test environments under heavy-load network conditions in Fig. 5. The greedy algorithm requires less running time and achieves better user demand satisfaction rates than RL, as it is a heuristic approach and proves to be optimal under heavy-load conditions.

### v3

@@ V-A Problem Transformation
We modify the original decision variable to enable the transformation of the original optimization problem into an ILP problem. Specifically, we define $\mathcal{M}^{q}$ as the collection of cooperative sets, each consisting of $q$ BSs selected from the BS set $\mathcal{M}$. For example, when there are three BSs in the network, i.e., =$\mathcal{M}=\{1,2,3\}$, the set of all combinations of two BSs selected from $\mathcal{M}$ is $\mathcal{M}^{2}=\{\{1,2\},\{1,3\},\{2,3\}\}$. Next, we introduce a binary variable $\delta_{\mathcal{B},n}$ to represent the selection of the cooperation set $\mathcal{B}$ for user $n$. When $\delta_{\mathcal{B},n}=1$, it indicates that cooperation set $\mathcal{B}\in\mathcal{M}^{q}$ is selected to serve user $n$, where $q\in\mathcal{M}$ and $|\mathcal{B}|=q$; if $\delta_{\mathcal{B},n}=0$, it means the cooperation set is not chosen. Let $o_{\mathcal{B},n}$ denote the number of subcarriers allocated to user $n$ by cooperation set $\mathcal{B}$. The constraint that the total number of allocated subcarriers does not exceed the available total number of subcarriers can be transformed into
@@
The objective function in (23) is expressed in terms of the transformed integer decision variables, while constraint (26) ensures that user demands are satisfied. The objective function and all constraints are linear, with decision variables constrained to integers, forming a standard ILP problem, denoted as $\mathcal{Z}_{0}$.
@@ V-B Two-Stage Resource Allocation Algorithm
To solve the aforementioned ILP problem, the TSRA algorithm is proposed. First, the feasibility pump algorithm quickly provides a high-quality feasible solution for subcarrier allocation. Then, the branch-and-cut method is applied to iteratively refine the solution to achieve optimality.
@@
$d(\bm{\chi},[\bm{\chi}^{\star}])=\sum_{[\delta_{j}^{\star}]=1}(1-\delta_{j})+\sum_{[\delta_{j}^{\star}]=0}\delta_{j}$ (29)
$+\sum_{[o_{j}^{\star}]=K}(K-o_{j})+\sum_{[o_{j}^{\star}]=0}o_{j}+\sum_{0<[o_{j}^{\star}]<K}(o_{j}^{+}+o_{j}^{-}),$
where $o_{j}^{+}$ and $o_{j}^{-}$ represent the non-negative differences between $o_{j}$ and $[o_{j}]$, corresponding to cases where $o_{j}>[o_{j}]$ and $o_{j}<[o_{j}]$, respectively. Then, the formulated LP problem in feasibility pump $\mathcal{Z}_{\text{LP}}(\bm{\chi},[\bm{\chi}^{\star}])$ can be expressed as follows with additional constraints on $o_{j}$:
@@
If $d(\bm{\chi},[\bm{\chi}^{\star}])=0$, a feasible solution to the original problem is obtained. The feasibility pump algorithm, forming the first stage of the proposed resource allocation method, is detailed in the initial part of Algorithm 1. Note that in Step 1 of Algorithm 1, $N_{\text{flip}}$ elements of $\bm{\chi}^{*}$ are replaced with $\bm{\chi}^{\star}$ rounded in the opposite direction to prevent an endless loop.
@@ 11                   $\bm{\chi}^{*}\leftarrow[\bm{\chi}^{\star}]$
14                   Select $N_{\text{flip}}$ variables of $\bm{\chi}^{*}$ with the highest $|\chi_{j}^{\star}-\chi_{j}^{*}|$ and replace $\chi_{j}^{*}$ with $\chi_{j}^{\star}$ rounded in the opposite direction
@@
17            Solve $\mathcal{Z}_{\text{LP}}(\bm{\chi}^{\star},\bm{\chi}^{*})$ of (30) to obtain $\bm{\chi}^{\star}$
@@ 44                   Add them to the LP relaxation of the problem $\mathcal{Z}_{y}$ and go back to Step 1
The latter part of Algorithm 1, starting from Step 1, outlines the branch-and-cut method. It utilizes a tree structure to represent the problem-solving process, where the root node corresponds to the LP relaxation of the original problem, providing an initial LB of the subcarrier energy cost. Similarly, the relaxed LP problem is solved by the interior-point method. The LB is the smallest energy cost from all explored LP relaxations, while the UB is the best energy cost from feasible integer solutions, initialized by the feasibility pump algorithm. Through branching, it partitions the feasible regions into smaller subspaces, creating subproblems as child nodes. At each node, cutting planes are introduced to further tighten the solution space, eliminating relaxed solutions that violate the integer constraints. Pruning is applied to discard branches where the LB exceeds the incumbent UB, reducing unnecessary searches.
By integrating the branch-and-cut with the feasibility pump, a feasible solution is quickly obtained and iteratively refined to optimality. The complexity of the proposed TSRA algorithm mainly stems from its second stage, which is the focus of the complexity analysis. Essentially, the branch-and-cut method establishes a search tree to explore and obtain the optimal solution. At each node, the LP relaxation of the problem is efficiently solved by the interior-point method. In the worst case, the time complexity is $\mathcal{O}(2^{|\mathcal{J}|})$ in terms of the number of nodes generated by a brute-force search. However, in practice, the number of actually explored nodes is generally much smaller due to effective pruning and cutting strategies. Moreover, the high-quality feasible solution from the first stage provides an initial UB that further facilitates pruning. Thus, the time complexity of the second stage is $\mathcal{O}(N_{en})$ in terms of the number of actually explored nodes $N_{en}$, where $N_{en}\ll 2^{|\mathcal{J}|}$. The space complexity is $\mathcal{O}(N_{an})$, where $N_{an}$ denotes the number of active nodes during the search, including those being processed and those branched out to be explored. Effective pruning typically ensures that $N_{an}<N_{en}$.
@@ VI-A Simulation Setup
In our simulation, BS is equipped with $N_{tx}=16$ transmit antennas, and the user is equipped with $N_{rx}=4$ receive antennas. Zero forcing (ZF) equalizer is employed on the user side. The center frequency is 3.5 GHz and the subcarrier spacing is 15 kHz, with a transmission time interval (TTI) of 1 ms and the numerology $\mu=0$. In a single TTI or time slot, the resource element (RE) grid for MIMO transmission consists of 2,016 REs, obtained from 144 subcarriers $\times$ 14 symbols, i.e., $K=144$. For simplicity, $OH$ is fixed at 0.14. The transmit power per subcarrier $p$ is set to 1 mW. In the greedy algorithm, $\eta_{\min}$ is fixed at 0.1. In TSRA, $N_{\text{flip}}$ is chosen to be 15 to avoid an endless loop and the large constant $\lambda$ is set to 1,000.
@@
Meanwhile, simulations are conducted based on the ray-tracing channel dataset: DeepMIMO [37]. Using accurate ray-tracing data obtained from Remcom Wireless InSite, clustered delay line (CDL) channels between BS and user are generated according to standard [38]. In this paper, we focus on using BS1,2,3 to transmit data to users in User Grid 1 of DeepMIMO’s O1 scenario, i.e., $M=3$. A total of 225 users are considered, spanning from R515 to R915 in O1, arranged in a 15-by-15 grid with 15 users evenly spaced along both the x-axis and the y-axis. Among them, 10 users are randomly selected in each case to have data rate demands, i.e., $N=10$. When generating the channel, the maximum number of channel paths is set to 10, though the actual number varies at different locations based on ray-tracing results.
User data demands at each time slot are known from the radio link control (RLC) buffer, calculated as the difference between total arrived and transmitted data. Within the relatively short simulation time, data arrivals per TTI for each user are assumed to be stable. The mean data arrival sizes vary across users to reflect differences in service requirements. Moreover, we simulate a snapshot of resource scheduling and thus assume that user locations remain unchanged.
VI-B Comparative Methods
We consider the practical implementation of scheduling algorithms under real-time conditions. Given the high computation efficiency of NNs and the widespread adoption of RL in addressing resource scheduling problems [27, 39, 40], the proposed algorithms are benchmarked against RL. Specifically, the dueling double deep Q-network (D3QN) algorithm is chosen as the RL method to solve the problem. Compared with DQN, D3QN enhances training stability, improves learning efficiency, and results in more accurate Q-value predictions [41, 42]. State in RL is defined as the concatenation of each user’s state, i.e., $\mathfrak{s}=(\mathfrak{s}_{1},\ldots,\mathfrak{s}_{N})$, where the state of user $n$ is expressed as $\mathfrak{s}_{n}=(\bm{loc}_{n},\vartheta_{n},R_{n}^{d},\hat{R}_{n})$. Since the problems differ under varying network conditions, the action and reward of RL should be designed accordingly. In heavy-load network conditions, action $\mathfrak{a}$ is designed to allocate one subcarrier from all DL-BSs to a selected user. Action mask is applied to filter available actions where the user’s demand is unmet, ensuring the generation of meaningful memories for RL to learn effectively [39]. The reward is the increase in the chosen user’s weighted demand satisfaction rate, i.e., $\mathfrak{r}=\vartheta_{\mathfrak{a}}$. The episode terminates when all $K_{\text{remaining}}$ subcarriers are allocated. In light-load network conditions, action is defined as allocating one subcarrier from a selected combination of cooperating BSs to a user. The reward is designed to guide the agent in meeting all user demands while minimizing subcarrier energy cost. If the highest SE TM is selected, user demands can be fully met, but at the cost of higher energy consumption. Conversely, if the most energy-efficient TM is used to serve users, it may result in unmet user demands after all subcarriers have been allocated. Therefore, the problem involves a tradeoff between SE and EE. To this end, we implement two types of RL to solve the problem, namely SE-prioritized and EE-prioritized RL. The reward is defined as the increase in the weighted sum of SE and EE resulting from the selected action $\mathfrak{a}$. It can be expressed as $\mathfrak{r}=\beta_{\text{SE}}\text{SE}_{\mathfrak{a}}+\beta_{\text{EE}}\text{EE}_{\mathfrak{a}}$, where $\beta_{\text{SE}}$ and $\beta_{\text{EE}}$ represent the importance of SE and EE, respectively. For SE-prioritized RL, the weights are set as $\beta_{\text{SE}}=5$ and $\beta_{\text{EE}}=1$. For EE-prioritized RL, the weights are set as $\beta_{\text{SE}}=1$ and $\beta_{\text{EE}}=5$. The episode ends when either all user demands are satisfied or all subcarriers are allocated.
VI-C Evaluation of Proposed Algorithms
To assess the performance of proposed algorithms under heavy-load and light-load network conditions, we generate independent time slot cases with random user locations, data arrivals, and priority weights. We randomly select 80% of the cases for RL’s training and use the remaining 20% for testing.
VI-C1 Evaluation of Proposed Greedy Algorithm
We first compare the proposed greedy algorithm with RL in heavy-load network conditions. Fig. 3 illustrates the running time and cumulative density function (CDF) of the weighted sum of user demand satisfaction rates $V$ under heavy-load network conditions. The greedy algorithm requires less running time and achieves a higher weighted sum of user demand satisfaction rates than RL, as it is a heuristic method and proves to be optimal under heavy-load conditions.